Close to 45% of smartphone users say they upgrade their phone every 2 years. In a random sample of 60 smartphone users, what is the probability that a. more than 35 smartphone users upgrade their phone every 2 years?
Answer #1

We have given n = 60 , x = 35 and p = 0.45

We have asked to find P ( More than 35 ) that is P ( x >35 )

P( x > 35 ) = P( x = 36 ) + P ( x = 37 ) + ....................... + P ( x = 58 ) + P ( x = 59 ) + P ( x = 60 )

By using Binomial distribution formula we get -

P( x > 35 ) = 0.0139 ( Rounded to 4 decimal places )

Hope this will help you. Thank you :)